This is from my son's homework. Would appreciate your help. Thanks in advance!
Explain how a(b + c) can be rewritten as (b + c)a and as ba + ca.
Specify which properties can be used.

Answer by richwmiller(17219)

Summary:
Commutative Laws: a + b = b + a
a * b = b * a
Associative Laws: (a + b) + c = a + (b + c)
(a * b) * c = a * (b * c)
Distributive Law: a * (b + c) = a * b + a * c